# Constrained free energy minimization for the design of thermal states and stabilizer thermodynamic systems

### v1

@@ VI Simulation results
Figure 4: The figure depicts the logarithm of the error metric in (122) versus the number of iterations, for the task of constrained energy minimization for the one-dimensional quantum Heisenberg model with nearest- and next-to-nearest-neighbor interactions and constraints on the total magnetizations in the $x$, $y$, and $z$ directions set to be 1, 0, and 1, respectively. The LMPW25 HQC algorithm converges in all cases (3, 5, and 6 qubits), with larger systems taking more iterations to converge.
@@ VII Conclusion
In this paper, we put forward several conceptual contributions, including framing constrained energy minimization as a method for designing ground states of Hamiltonians and constrained free energy minimization as a method for designing thermal states of Hamiltonians. We argued how the LMPW25 HQC algorithms can be used for designing these Hamiltonians and how this application has implications for designing molecules and materials with desired properties. We also introduced stabilizer thermodynamic systems as thermodynamic systems for which the Hamiltonian is constructed from the stabilizer operators of a stabilizer code and the conserved, non-commuting charges are constructed from the code’s logical operators. Then we demonstrated how the LMPW25 HQC algorithms can be understood as alternative methods for encoding quantum information into quantum error-correcting codes, which could be a useful approach whenever low-temperature thermal state preparation methods are experimentally accessible.
@@
The numerical contribution of our paper involved simulating the LMPW25 algorithms for several key thermodynamic systems of interest, including the one- and two-dimensional quantum Heisenberg models with nearest-neighbor and next-to-nearest-neighbor interactions and several stabilizer thermodynamic systems (the one-to-three qubit repetition code, the two-to-four qubit error-detecting code, and the perfect one-to-five qubit code). To test the performance of the LMPW25 algorithms on all of these examples, we employed standard SDP solvers to determine the true values of the constrained energy minimization problems. The LMPW25 classical algorithms can be understood as establishing alternative benchmarks for the HQC algorithms: the former algorithms act as noiseless simulators of the HQC algorithms, which focus exclusively on converging when measurement shot noise is not present, while our simulations of the latter incorporate measurement shot noise.
@@
Going forward from here, there are several avenues to pursue for future research. First, we think it would be worthwhile to employ matrix product states for simulating the thermal states needed in the LMPW25 classical algorithms, as used in prior work [95, 67, 96]. Very likely such an approach would lead an advanced classical algorithm and serve several purposes: first, it could be a useful algorithm on its own for understanding quantum thermodynamic systems; second, it could be helpful for solving or approximating the design of ground and thermal states of controllable Hamiltonians; third, it could be useful as a benchmark to set against the LMPW25 HQC algorithms. Regarding the latter point, any claimed quantum speedup for the LMPW25 HQC algorithms would have to outperform any possible classical algorithm for this task, and it is likely that approaches involving matrix product states would lead to the fastest possible classical algorithms.
@@
We also left it open to test the performance of the LMPW25 HQC algorithms on currently available quantum computers. In doing so, it would be necessary to incorporate the techniques of quantum error correction and mitigation in order to limit the effects of noise, and one would also need to incorporate various advanced methods for thermal state preparation [29, 30, 31, 32, 33, 34, 35, 36] in order to have efficient approaches for this necessary step. As discussed in [16], in order to have an efficient HQC algorithm, it would be necessary to find a quantum thermodynamic system (Hamiltonian and conserved charges) for which one can efficiently prepare low-temperature thermal states.
@@ Acknowledgements.
We are grateful to Nicole Yunger Halpern for a helpful discussion on quantum thermodynamics in the presence of conserved, non-commuting charges. We also thank Zoë Holmes, Peter McMahon, Karan Mehta, Erich Mueller, and Darren Pereira for helpful discussions. MM acknowledges support from the Swiss National Science Foundation [grant number 200021-219329]. MC and JK acknowledge support from the Cornell Engineering Learning Initiative program. IL, ML, and KW are grateful to the School of Electrical and Computer Engineering for hospitality during their summer internships. NL acknowledges funding from the Science and Technology Commission of Shanghai Municipality (STCSM) grant no. 24LZ1401200 (21JC1402900), NSFC grants no. 12471411 and no. 12341104, the Shanghai Jiao Tong University 2030 Initiative, and the Fundamental Research Funds for the Central Universities. MMW acknowledges support from the National Science Foundation under grant no. 2329662 and from the Cornell School of Electrical and Computer Engineering.
@@
ML: Data curation, Resources, Software, Validation, Visualization, Writing - review & editing.
SR: Software, Supervision, Validation, Writing - review & editing.
@@ Appendix A Closeness of low-temperature thermal states and ground states of Hamiltonians
By Propositions 911, and 13, we have that
@@ Appendix F Proof of Theorem 8
$U\overline{\sigma}_{i_{1}}^{{\dagger}}\cdots\overline{\sigma}_{i_{k}}U$ $=I^{\otimes n-k}\otimes\sigma_{i_{1}}\otimes\cdots\otimes\sigma_{i_{k}}.$ (300)
@@ G.1 One-dimensional quantum Heisenberg model with nearest-neighbor interactions
Figure 5: The figure depicts the logarithm of the error metric in (122) versus the number of iterations, for the first-order classical and HQC algorithms and the second-order classical algorithm for the task of constrained energy minimization for the five-qubit, one-dimensional quantum Heisenberg model with nearest-neighbor interactions and constraints on the total magnetizations in the $x$, $y$, and $z$ directions set to be 1, 0, and 1, respectively. All of the algorithms converge, but the HQC algorithm requires more iterations to converge due to sampling noise inherent in it.
@@ G.2 Two-dimensional quantum Heisenberg model with nearest-neighbor interactions
Figure 6: The figure depicts the logarithm of the error metric in (122) versus the logarithm of the number of iterations, for the first-order classical and HQC LMPW25 algorithms for the task of constrained energy minimization for the six-qubit, two-dimensional quantum Heisenberg model with nearest-neighbor interactions and constraints on the total magnetizations in the $x$, $y$, and $z$ directions set to be 1, 0, and 1, respectively. Both algorithms converge, but the HQC algorithm requires more iterations to converge due to sampling noise inherent in it.
Figure 6 depicts the performance of the first-order classical and HQC LMPW25 algorithms for the six-qubit, two-dimensional quantum Heisenberg model with nearest-neighbor interactions and constraints on the total magnetizations in the $x$, $y$, and $z$ directions set to be 1, 0, and 1, respectively.
@@ G.3 One-to-three-qubit repetition code
Figure 7 plots the performance of the LMPW25 classical first-order algorithm with and without a warm start, demonstrating that warm starting the algorithm requires fewer iterations for convergence to occur. Figure 8 is a similar plot, but using instead the LMPW25 HQC first-order algorithm, demonstrating a significant performance improvement by warm starting.
Figure 7: The figure depicts the logarithm of the error metric in (122) versus the number of iterations, for the LMPW25 classical first-order algorithm for the task of constrained energy minimization for the one-to-three-qubit repetition code. In this case, the plot compares the performance with and without a warm start for the algorithm, as discussed in Section V.1.
Figure 8: The figure depicts the logarithm of the error metric in (122) versus the number of iterations, for the LMPW25 HQC first-order algorithm for the task of constrained energy minimization for the one-to-three-qubit repetition code. In this case, the plot compares the performance with and without a warm start for the algorithm, as discussed in Section V.1.
@@ G.4 Two-to-four-qubit quantum error-detecting code
Here, we test the performance of the LMPW25 algorithms for encoding a Bell state $|\Phi^{+}\rangle\!\langle\Phi^{+}|$, where $|\Phi^{+}\rangle\coloneqq\frac{1}{\sqrt{2}}\left(|00\rangle+|11\rangle\right)$, into the four-qubit error-detecting code. The representation of this state in the Pauli basis is as follows:
@@
These constraints are chosen in accordance with the representation in (324), in order to ensure that the encoded two-qubit state is indeed a Bell state. The goal is to compare the performance of all of the LMPW25 algorithms for the constrained energy minimization problem. Simulations results for this example will appear in an updated version of our paper.

### v2

@@ VI Simulation results
Figure 5: The figure depicts the average of the logarithm of the error metric in (122) over five independent runs, plotted against the number of iterations, with shaded regions indicating one standard deviation. The task is constrained energy minimization for the one-dimensional quantum Heisenberg model with nearest- and next-to-nearest-neighbor interactions, and constraints on the total magnetizations in the $x$, $y$, and $z$ directions set to be 1, 0, and 1, respectively. The LMPW HQC algorithm converges in all cases (3, 4, 5, and 6 qubits), with larger systems taking more iterations to converge.
Our final set of simulations involves the one-dimensional quantum Heisenberg model with nearest- and next-to-nearest-neighbor interactions and with the charges set to the total magnetizations in the $x$, $y$, and $z$ directions. The goal is to solve the constrained energy minimization problem reviewed in Section II.2.1, with the expectation values on the total magnetizations in the $x$, $y$, and $z$ directions set to be 1, 0, and 1, respectively. Figure 4 demonstrates the performance of the first-order LMPW classical algorithm as the number of qubits in the model increases (specifically, the figure plots results for three, four, five, and six qubits). The figure plots the error metric in (122) versus the number of iterations, demonstrating that every case converges to the true value as the number of iterations increases.
Figure 5 demonstrates the performance of the first-order HQC algorithm for the same setup. The algorithm converges in all cases, but requires more iterations than the classical algorithm and exhibits noticeable noise in proximity to the optimal value. As expected, this behavior arises from the sampling noise inherent in the HQC algorithm. To mitigate the increased computational cost and instability associated with this noise, we employed a looser convergence threshold on the gradient norm, which also accounts for the slightly higher final error metric reached by the HQC algorithm compared to its classical counterpart. The number of shots used per iteration was set according to [16, Eq. (F17)], with $\varepsilon=10^{-4}$ and $\delta=10^{-2}$. In this case, the norm of the Pauli coefficients of the total magnetizations is $n$, where $n$ is the number of qubits.
VI.1 Comparison of Standard Gradient Ascent and Nesterov’s Accelerated Method
Figure 6: The figure depicts the logarithm of the error metric in (122) versus the number of iterations for the LMPW classical first-order algorithm, comparing the standard gradient ascent method with Nesterov’s accelerated version. The task is constrained energy minimization for the two-dimensional, four-qubit quantum Heisenberg model with nearest- and next-to-nearest-neighbor interactions and constraints on the total magnetizations in the $x$, $y$, and $z$ directions set to be 1, 0, and 1, respectively.
Figure 7: The figure depicts the average of the logarithm of the error metric in (122) over five independent runs, plotted against the number of iterations, with shaded regions indicating one standard deviation, for the LMPW HQC first-order algorithm, comparing the standard gradient ascent method with Nesterov’s accelerated version. The task is constrained energy minimization for the two-dimensional, four-qubit quantum Heisenberg model with nearest- and next-to-nearest-neighbor interactions and constraints on the total magnetizations in the $x$, $y$, and $z$ directions set to be 1, 0, and 1, respectively.
Throughout this work, we employed the standard gradient ascent algorithm for both the classical and HQC implementations of the first-order optimization procedure. However, several well-known modifications of the standard gradient ascent method can lead to improved convergence performance, especially in concave settings. One such modification is Nesterov’s accelerated gradient method [81], which introduces a momentum term to anticipate the direction of the next update, resulting in improved convergence speed.
While the standard gradient ascent update rule, given in (19), directly updates the parameters based on the gradient evaluated at the current iterate, Nesterov’s accelerated gradient ascent maintains an auxiliary sequence $v_{m}$, which incorporates a momentum term, and performs the update as follows:
$v_{m+1}$ $=\tau v_{m}+\eta\nabla f(\mu_{m}+\tau v_{m}),$ (136)
$\mu_{m+1}$ $=\mu_{m}+v_{m+1},$ (137)
where $\tau\in[0,1)$ is the momentum parameter and $\eta>0$ is the learning rate. The key idea is to compute the gradient not at the current iterate $\mu_{m}$, but at the projected position $\mu_{m}+\tau v_{m}$, effectively introducing a look-ahead mechanism that often leads to faster convergence.
Figure 6 illustrates the performance of the standard and Nesterov-accelerated gradient ascent algorithms for the classical first-order optimization of the two-dimensional four-qubit Heisenberg model with nearest- and next-to-nearest-neighbor interactions. While both methods converge to the same solution, the Nesterov-accelerated method does so in significantly fewer iterations. Figure 7 presents the corresponding results for the HQC first-order optimization applied to the same task, again demonstrating that the Nesterov-accelerated method converges in significantly fewer iterations than the standard gradient ascent method.
@@ VII Conclusion
In this paper, we put forward several conceptual contributions, including framing constrained energy minimization as a method for designing ground states of Hamiltonians and constrained free energy minimization as a method for designing thermal states of Hamiltonians. We argued how the LMPW HQC algorithms can be used for designing these Hamiltonians and how this application has implications for designing molecules and materials with desired properties. We also introduced stabilizer thermodynamic systems as thermodynamic systems for which the Hamiltonian is constructed from the stabilizer operators of a stabilizer code and the conserved, non-commuting charges are constructed from the code’s logical operators. Then we demonstrated how the LMPW HQC algorithms can be understood as alternative methods for encoding quantum information into quantum error-correcting codes, which could be a useful approach whenever low-temperature thermal state preparation methods are experimentally accessible.
@@
The numerical contribution of our paper involved simulating the LMPW algorithms for several key thermodynamic systems of interest, including the one- and two-dimensional quantum Heisenberg models with nearest-neighbor and next-to-nearest-neighbor interactions and several stabilizer thermodynamic systems (the one-to-three qubit repetition code, the two-to-four qubit error-detecting code, and the perfect one-to-five qubit code). To test the performance of the LMPW algorithms on all of these examples, we employed standard SDP solvers to determine the true values of the constrained energy minimization problems. The LMPW classical algorithms can be understood as establishing alternative benchmarks for the HQC algorithms: the former algorithms act as noiseless simulators of the HQC algorithms, which focus exclusively on converging when measurement shot noise is not present, while our simulations of the latter incorporate measurement shot noise.
@@
Going forward from here, there are several avenues to pursue for future research. First, we think it would be worthwhile to employ matrix product states for simulating the thermal states needed in the LMPW classical algorithms, as used in prior work [96, 67, 97]. Very likely such an approach would lead an advanced classical algorithm and serve several purposes: first, it could be a useful algorithm on its own for understanding quantum thermodynamic systems; second, it could be helpful for solving or approximating the design of ground and thermal states of controllable Hamiltonians; third, it could be useful as a benchmark to set against the LMPW HQC algorithms. Regarding the latter point, any claimed quantum speedup for the LMPW HQC algorithms would have to outperform any possible classical algorithm for this task, and it is likely that approaches involving matrix product states would lead to the fastest possible classical algorithms.
@@
We also left it open to test the performance of the LMPW HQC algorithms on currently available quantum computers. In doing so, it would be necessary to incorporate the techniques of quantum error correction and mitigation in order to limit the effects of noise, and one would also need to incorporate various advanced methods for thermal state preparation [29, 30, 31, 32, 33, 34, 35, 36] in order to have efficient approaches for this necessary step. As discussed in [16], in order to have an efficient HQC algorithm, it would be necessary to find a quantum thermodynamic system (Hamiltonian and conserved charges) for which one can efficiently prepare low-temperature thermal states.
@@ Acknowledgements.
We are grateful to Nicole Yunger Halpern for a helpful discussion on quantum thermodynamics in the presence of conserved, non-commuting charges. We also thank Zoë Holmes, Peter McMahon, Karan Mehta, Erich Mueller, and Darren Pereira for helpful discussions. MM acknowledges support of the NCCR MARVEL, a National Centre of Competence in Research, funded by the Swiss National Science Foundation (grant number 205602). MC and JK acknowledge support from the Cornell Engineering Learning Initiative program. IL, ML, and KW are grateful to the Cornell School of Electrical and Computer Engineering for hospitality during their summer internships. NL acknowledges funding from the Science and Technology Commission of Shanghai Municipality (STCSM) grant no. 24LZ1401200 (21JC1402900), NSFC grants no. 12471411 and no. 12341104, the Shanghai Jiao Tong University 2030 Initiative, and the Fundamental Research Funds for the Central Universities. MMW acknowledges support from the National Science Foundation under grant no. 2329662 and from the Cornell School of Electrical and Computer Engineering.
@@
ML: Data curation, Resources, Software, Validation, Visualization.
SR: Software, Supervision, Validation.
@@ Appendix A Closeness of low-temperature thermal states and ground states of Hamiltonians
By Propositions 9, 11, and 13, we have that
@@ Appendix F Proof of Theorem 8
$U^{{\dagger}}\overline{\sigma}_{i_{1}}\cdots\overline{\sigma}_{i_{k}}U$ $=I^{\otimes n-k}\otimes\sigma_{i_{1}}\otimes\cdots\otimes\sigma_{i_{k}}.$ (302)
@@ Appendix G Other simulation results
The technical details of the algorithmic implementations and simulation setup used for the results presented in this appendix are similar to those described in Section VI of the main text. We therefore refer the reader to that section for explanations of implementation-specific aspects of the classical and HQC algorithms.
@@ G.1 One-dimensional quantum Heisenberg model with nearest-neighbor interactions
Figure 8: The figure depicts the logarithm of the error metric in (122) versus the number of iterations, for the first-order classical and HQC algorithms and the second-order classical algorithm for the task of constrained energy minimization for the five-qubit, one-dimensional quantum Heisenberg model with nearest-neighbor interactions and constraints on the total magnetizations in the $x$, $y$, and $z$ directions set to be 1, 0, and 1, respectively. All of the algorithms converge, but the HQC algorithms, shown as the average over five independent runs with shaded regions denoting one standard deviation, require more iterations to converge due to sampling noise inherent in them.
@@ G.2 Two-dimensional quantum Heisenberg model with nearest-neighbor interactions
Figure 9: The figure depicts the logarithm of the error metric in (122) versus the logarithm of the number of iterations, for the first-order classical and HQC LMPW algorithms for the task of constrained energy minimization for the six-qubit, two-dimensional quantum Heisenberg model with nearest-neighbor interactions and constraints on the total magnetizations in the $x$, $y$, and $z$ directions set to be 1, 0, and 1, respectively. Both algorithms converge, but the HQC algorithm, shown as the average over five independent runs with the shaded region denoting one standard deviation, requires more iterations to converge due to sampling noise inherent in it.
Figure 9 depicts the performance of the first-order classical and HQC algorithms and the second-order classical algorithm for the six-qubit, two-dimensional quantum Heisenberg model with nearest-neighbor interactions and constraints on the total magnetizations in the $x$, $y$, and $z$ directions set to be 1, 0, and 1, respectively. Given the substantial computational cost associated with Hessian estimation in the HQC setting, we did not include the second-order HQC algorithm in this simulation.
@@ G.3 One-to-three-qubit repetition code
Figure 10 plots the performance of the LMPW classical first-order algorithm with and without a warm start, showing that warm starting leads to immediate convergence in just one iteration. Figure 11 presents a similar plot for the LMPW HQC first-order algorithm, which likewise benefits from a warm start by converging in a single iteration.
Figure 10: The figure depicts the logarithm of the error metric in (122) versus the number of iterations, for the LMPW classical first-order algorithm for the task of constrained energy minimization for the one-to-three-qubit repetition code. In this case, the plot compares the performance with and without a warm start for the algorithm, as discussed in Section V.1. The warm-started algorithm converges in a single iteration.
Figure 11: The figure depicts the average of the logarithm of the error metric in (122) over five independent runs, plotted against the number of iterations, with shaded regions indicating one standard deviation, for the LMPW HQC first-order algorithm for the task of constrained energy minimization for the one-to-three-qubit repetition code. In this case, the plot compares the performance with and without a warm start for the algorithm, as discussed in Section V.1. The warm-started algorithm converges in a single iteration, and the error bars reflect the standard deviation across the five runs.
@@ G.4 Two-to-four-qubit quantum error-detecting code
Here, we test the performance of the LMPW algorithms for encoding a Bell state $|\Phi^{+}\rangle\!\langle\Phi^{+}|$, where $|\Phi^{+}\rangle\coloneqq\frac{1}{\sqrt{2}}\left(|00\rangle+|11\rangle\right)$, into the four-qubit error-detecting code. The representation of this state in the Pauli basis is as follows:
@@
These constraints are chosen in accordance with the representation in (326), in order to ensure that the encoded two-qubit state is indeed a Bell state. Figure 12 plots the performance of the LMPW algorithms for this constrained energy minimization problem.
Figure 12: The figure depicts the logarithm of the error metric in (122) versus the number of iterations, for the first-order classical and HQC algorithms and the second-order classical algorithm for the task of constrained energy minimization for the two-to-four-qubit quantum error-detecting code. All of the algorithms converge, but the HQC algorithms, shown as the average over five independent runs with the shaded region denoting one standard deviation, require more iterations to converge due to sampling noise inherent in them.